20 g of ice in 10 g copper calorimeter is originally at -30 degree C. how much steam at 100 degree C must be condensed in the calorimeter if the ice is to be changed to water and heated to 40 degree C

Grade:11

1 Answers

Eshan
5 years ago
Total water equivalent mass of the system is20g+10g=30g

Hence the amount of heat needed to reach its temperature to40^{\circ}C=(30)(0.5)(0-(-30))+(30)(80)+(30)(1)(40-0)
Hence this should be equal to the amount of heat given out by condensingmmass of stem.
\implies m(540)=(30)(0.5)(0-(-30))+(30)(80)+(30)(1)(40-0)
\implies m=7.5g